OpenStudy (logan13):

What is the percent of increase from 50 to 90? A. 40% B. 44% C. 56% D. 80%

OpenStudy (logan13):

@EclipsedStar plz i beg hlp me plz

OpenStudy (mathstudent55):

\(percent~ change = \dfrac{new ~number - old ~number}{old~number} \times 100\)

OpenStudy (logan13):

math student i have 4 more

OpenStudy (mathstudent55):

In your case, the new number is 90 and the old number is 50.

OpenStudy (mathstudent55):

Ok, let's do them one at a time. Use the formula above and plug in 90 and 50 and figure it out. Let me know what you get.

OpenStudy (logan13):

40 %

OpenStudy (logan13):

@EclipsedStar = thank u anyways

OpenStudy (mathstudent55):

That's not correct. Can you show me your numbers in the formula?

OpenStudy (logan13):

okay i suck um like how i did it

OpenStudy (mathstudent55):

Did you try it at all or did you just pick the first choice at random?

OpenStudy (mathstudent55):

These are the numbers you should use in the formula: \(percent~ change = \dfrac{90 - 50}{50} \times 100\)

OpenStudy (mathstudent55):

What is 90 - 50 = ?

OpenStudy (logan13):

50%

OpenStudy (logan13):

40

OpenStudy (mathstudent55):

Good. Use a calculator now: what is 40/50 = ?

OpenStudy (logan13):

80

OpenStudy (mathstudent55):

40/50 = 0.8 Then 0.8 * 100 = 80 The answer is 80% increase